A student on a piano stool rotates freely with an angular speed of 3.07 rev/s. The student holds a 1.27-kg mass in each outstretched arm, 0.766 m from the axis of rotation. The combined moment of inertia of the student and the stool, ignoring the two masses, is 5.50 kg*m2, a value that remains constant. As the student pulls his arms inward, his angular speed increases to 3.40 rev/s. How far are the masses from the axis of rotation at this time, considering the masses to be points?

Calculate the initial kinetic energy of the system.

Calculate the final kinetic energy of the system.
Answer #1

using conservation of angular momentum

( 2 m r^2 + I) w1 = (2 m d^2 + I) w

( 2* 1.25* 0.766^2 + 5.5)* 3.07 = ( 2* 1.25* d^2 + 5.5)* 3.4

d = 0.5624 m

======

b)

initial kE

kEi = 0.5 I w1^2 = 0.5* (5.5 + 2* 1.25* 0.766^2)* (3.07* 2pi) ^2

kEi = 1296.12 J

final kE

kEf = 0.5* (2* 1.25* 0.5624^2 + 5.5)* ( 3.4* 2pi)^2

kEf = 1435.45 J

======

Comment before rate in case any doubt, will reply for sure.. goodluck